OpenStudy (anonymous):

has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint? Im not sure how to solve this

OpenStudy (lgbasallote):

hint: \[\frac{x_1 + x_2}{2} = m_x\] you are given x_1 as 25 and m_x as 18 so solve for x_2 does that help?

OpenStudy (lgbasallote):

btw m_x means x coordinate of the midpoint

OpenStudy (anonymous):

oh okay thanx!

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

so it would be 11?

OpenStudy (lgbasallote):

\[\frac{25 + x_2}{2} = 18\] \[25 + x_2 = 36\] \[x_2 = 36 - 25\] \[x_2 = 11\] you are right :)

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (lgbasallote):

welcome ^_^